To:

From: Sung-Hoon Ko, Fax; (315)476-9206 Date: 8/8/95

I want to purchase home theater systems. Would you please tell me your best price of the following items. Thank you very much. (Please just send this paper back after filling out price. If you don’t have items just write ’no’ )

TV set : Toshiba CF35D90, price: Toshiba CF35D70, price:

VCR : SONY SLV-920HF, price:

RECEIVER : ONKYO TX-SV727PRO, price:

SPEAKERS : Definitive Technology BP-10, price: Atlantic Technology 253C(center speaker), price: Atlantic Technology 254SR(surround speakers), price(1 pair):

LD player : Pioneer CLD-M403, price: Pioneer CLD-V850, price: SAMSUNG DV730K, price: SONY MDP-600, price: YAMAHA CDV-W901, price:

OMNI Mount RST, price:

Shippin and Handling Charge to Syracuse NY, price:

Any Suggestion: